# The Peripatetic Hater: Predicting Movement Among Hate Subreddits

### v1

@@ Methods
To understand whether users who initially post in hate subreddits are at risk of becoming engaged in hate subreddits of other types, we employ a process to match each user from a given hate subreddit (a “joiner”) with a similar “counterpart” user who did not engage in a hate subreddit before the joiner, as described in Schmitz et al. 2024. For a given hate subreddit, we only consider users who have posted in the subreddit as their first hate subreddit. For each hate subreddit, the process consists of finding subreddits that users from the hate subreddit are most active in relative to their total size, and then sampling potential counterparts from those subreddits. Each user from the hate subreddit is matched to a counterpart using Mahalanobis distance matching. The features used to measure the Mahalanobis distance between users are each user’s total karma, their total number of submissions, their total number of comments, their account creation date, and the total number of posts they made in the top 50 subreddits in which the members of the given hate subreddit were most active. For each counterpart user, all features were measured prior to the month the treatment user posted in the hate subreddit. From these sets of users, we calculate the mean number of subreddits of a different type than the given hate subreddit that are joined within our chosen time threshold of six weeks following the users’ first post in the given subreddit. We report the average values for each category of subreddit as a whole. As counterpart users by definition do not participate in the hateful subreddit of their treatment counterpart, the number of subreddits joined by each counterpart user is measured within the same time period as the hate subreddit user to whom they are matched.
In addition to quantifying the level of movement across subreddits through sets of matched users, for each hate subreddit category, we measure the proportion of users who originate in that category that eventually also post in a hate subreddit belonging to another category of hate.
Evaluating peripatetic users’ use of language associated with destination subreddits in their original subreddit
@@
Using the lexicon of ingroup language associated with each category of hate subreddit, we seek to understand how users who post in multiple categories of hate subreddits introduce such language to different types of communities. For each user, we identify the category of hate subreddit in which they first posted. Then, in each category, we identify users who eventually posted in a different category of hate subreddit. We then compare the language employed by such users to the language used by those who exclusively posted within a single category. For peripatetic users, we only considered posts prior to posting in other hate subreddits, and all posts were collected within each user’s first three days of activity in any hate subreddit. For each combination of origin and destination subreddits, we count the number of terms used from the lexicon of the destination category relative to the number of terms used from all other categories. These are counted separately for peripatetic users and non-peripatetic users. By counting the number of terms associated with each category, we are able to calculate the odds that a peripatetic user employs terms from their destination category within the original category before they posted in the destination category, relative to the odds that non-peripatetic users would use the same terms (i.e., the odds ratio). We assess the statistical significance of changes in language use using Fisher’s exact test.
In addition to comparing the ingroup language used by peripatetic users vs. non-peripatetic users, we also measure the relative increase in ingroup language use from an alternate category after peripatetic users become active in said alternate category, as well as the percentage change in their activity rate (number of posts per week) in all hate subreddits. For each peripatetic user, a symmetric window of time is used to measure behavior before and after becoming peripatetic (e.g., if a user posts in a misogynistic subreddit three weeks after posting in a racist subreddit, their behavior in that three weeks is measured as the “before” period and their behavior three weeks after posting in the misogynistic subreddit is measured as the “after” period).
@@
In addition to understanding how content associated with hate subreddits in which peripatetic users eventually post appears in their postings in their initial community, we are also interested in the general themes expressed by peripatetic users, unbounded by a specific source. To understand these themes, we employ topic modeling using BERTopic (Grootendorst 2022). BERTopic uses pre-trained transformer models to embed text, then clusters those embeddings to form topics. The topics are then summarized using the highest TF-IDF words for each cluster. For each hate subreddit, we fit a topic model using the first three days of posts for each user. For peripatetic users who posted in another subreddit category within the first three days, only posts prior to this event are considered. As a preprocessing step, we remove URLs and the strings “[deleted]” and “[removed]” from all posts. We use the “outlier reduction” feature of BERTopic to assign posts that were not assigned to a topic when initially fitting the model to the closest topic based on the cosine similarity of the text embeddings. We then merge the topic models of all subreddits of the same category. After merging the topic models, we count the number of posts in the top 100 topics that peripatetic users and non-peripatetic users employ. For each topic, the odds of a peripatetic user making a post that fits within the given topic are calculated relative to the odds of a non-peripatetic user making a post on the same topic. When comparing the odds ratios of topics, only topics that appear in at least ten percent of the subreddits within the given category are considered, as some topics may come from individual subreddits that are more or less likely to contain peripatetic users. The five topics with the highest and lowest log-odds ratios for each category are reported.
@@
To probe the dynamics wherein users join additional categories of hate subreddits, we initially seek to establish whether certain factors are predictive of a user joining a different category of hate subreddit in the first place. Specifically, as we establish how language spreads among the categories of hate subreddits, we are interested in how the text to which users are exposed, and the text that they themselves employ, is indicative of their participation in each type of hate subreddit. To examine the text to which users are exposed, we utilize the posts to which the user replies, as we are confident that the user has read these posts. However, it is important to note that this approach a) excludes a great deal of text that the users likely read within the given subreddit, and b) may reflect self-selection, as the text to which users reply is likely to be especially related to their interests. We frame this as a multi-label classification problem, where we build a neural network trained on the embeddings of each type of text, and have the neural network predict whether a user subsequently joins a subreddit in each category. The text each user writes or replies to in any subreddit within three days of posting in their initial hate subreddit (including their first post in that subreddit) is embedded and fed into our model as input features (we explain embedding below). To prevent data leakage, if, within the three-day window, a user posts in a hate subreddit outside of the category in which they first posted, only posts before that point are considered. As we are primarily interested in users posting in other categories of subreddits, and not continuing to post in their original category, we do not consider a user making multiple posts in their original category as a positive example. Using this strategy, we compile a dataset of approximately 100k randomly sampled users, 70% of which was used for training, with 15% being used each for validation and testing.
@@
For each user in the dataset, all of the text that they post in their initial subreddit category is chronologically concatenated into the same document, with each post separated by the ‘$</s>$’ token. The text to which the users reply is separately concatenated in the same way. Each document is embedded into a 768-dimensional vector using Longformer (Beltagy, Peters, and Cohan 2020). The vectors are then concatenated, resulting in a 1,538-dimensional input layer for the neural network. As some data are missing from Pushshift (Gaffney and Matias 2018), there are some cases in which all of the posts to which a user replied could not be retrieved; in those cases, a single token (‘UNK’) was used as a placeholder for the parent text. The category of subreddit in which each user initially posted, as well as the categories the users to whom they replied posted prior to their interaction, are input as a one-hot encoding vector into the neural network at the second layer. The final neural network is three layers, with 774 dimensions in the second layer (where six of the input features derive from the one-hot encoding describing the initial subreddit and the subreddits of the parent users) and 384 dimensions in the third layer (i.e., each intermediate layer divides the number of features from the text vectors by a factor of two, with additional features in the second layer due to the one-hot encoding). At each hidden layer, a dropout of 0.6 is used, as well as the leaky ReLU activation function. The softmax activation function is used at the final output layer. We perform the training on an internal computing cluster using an NVIDIA RTX A6000 GPU.
@@
Finally, as there is variation in the performance of the model among individual runs, we take random splits of training, validation, and testing and train the model 50 times to calculate the mean and standard errors of the performance metrics. In addition, we only report performance for each category considering users who did not originate in that category. In contrast, the performance would be trivially almost perfect (ROC-AUC$>0.95$) if we also predict whether a user will remain in their present category, as that occurs for the vast majority of users.
@@ Results
Figure 2 shows the average number of posts predicted as hate speech for each cluster identified by the $k$-means algorithm. From the distributions, it is clear that one cluster exhibits a high degree of xenophobia and racism, another cluster exhibits high homophobia and transphobia, and the last cluster is high in misogyny. Therefore, we refer to the categories of subreddits as “racist,” “anti-LGBTQ,” and “misogynistic,” respectively. We find 69 racist subreddits, 32 anti-LGBTQ subreddits, and 19 misogynistic subreddits. Figure 1 shows the distributions of average hate scores for each subreddit plotted in two-dimensional space. Observing the plot, the categories form distinct clusters. Our results are robust to different clustering methods.
@@
Figure 3: Proportion of users who post in alternate subreddits within six weeks days after becoming active in their initial subreddit. Subreddits of the same category as the user’s initial category are not counted. Error bars represent standard errors.
@@
Figure 3 displays the proportion of newcomers who post in hate subreddits with matched treatment/counterpart pairs. It is apparent that, on average, the users who post in each hate subreddit subsequently participate in many more subreddits of another type than do users who do not post in the given hate subreddit (6-8 times higher depending on the type of subreddit; unweighted mean across all three categories: 6.9). The proportion of users who originate in racist subreddits is lower than the number joined by those who originate in the other categories, likely because there are more subreddits in the racist category. Using the Wilcoxon signed-rank test to assess differences between treatment and counterpart groups for each category, we find that the observed results are extremely unlikely to happen by chance, with $p<10^{-5}$ for each result.
RQ1: The probability of posting in an alternate category of hate subreddit does not depend on the original category.
@@
Figure 4: Heatmaps representing peripatetic user behavior (A) Proportion of users from the original subreddit type that subsequently posted in the alternate subreddit. (B) Odds ratios of peripatetic vs. non-peripatetic users using language from the alternate hate subreddit lexicon within the origin subreddit type. (C) Change in use of language from alternate hate subreddit lexicons after posting in alternate hate subreddits. (D) Percentage changes in activity rates after posting in alternate hate subreddits. Diagonals in A, C, and D represent users who do not post in alternate subreddits but post in multiple subreddits of the same category. Stars represent statistical significance ($p<0.001$)
Figure 4A displays the rate at which members who originate in each subreddit category subsequently post in other categories of subreddits. From the figure, it is immediately apparent that, for each subreddit type, the probability of ending up there as an alternate category is the same regardless of the original category. The average proportions in the columns of the heatmap (excluding cells in which the alternate and original categories are the same) are $5.5\%,0.8\%$, and $2.2\%$, which, when normalized, are almost identical to the distribution of total users that post in each category ($65.7\%$, $9.3\%$, and $25.8\%$).
@@
RQ2: Peripatetic users introduce language from alternate hate subreddits to the hate subreddit category in which they initially posted.
@@
Figure 4B displays the odds of peripatetic users posting material, in the category to which they first contributed, that employs language common in the peripatetic user’s subsequent subreddit category, relative to the odds that non-peripatetic users employ such language in the first category. The heatmap reveals that every result is statistically significant and is associated with peripatetic users employing more ingroup language from the destination subreddit relative to non-peripatetic users. The most prominent differences occur when the destination category is racist and the original category is misogynistic, and when the destination is misogynistic and the origin is anti-LGBTQ. Conversely, peripatetic users who eventually join misogynistic subreddits after being active in a racist subreddit exhibit a less substantial increase in misogynistic ingroup language within racist subreddits, though the result is still statistically significant.
@@
Figure 5 displays the ten most common topics within each category of subreddit. Among these topics, it is apparent that types of hate other than those that form the basis for the categorization of the subreddits are discussed. For example, topics represented by the word “race,” “Muslim,” or antisemitic slurs appear in the racist category, and topics represented by the word “gay” appear in the anti-LGBTQ category. In the misogynistic category, discussions about the gender pay gap and sexual violence are among the most frequent.
Figure 6: Topics with the five highest/lowest log-odds ratios of use between peripatetic users and non-peripatetic users. Positive values indicate topics that are over-represented among peripatetic users. Topics with an adjusted p-value greater than 0.01 according to Fisher’s exact test are lighter to show statistical insignificance. Stopwords were removed from topic representations.
The topics that peripatetic users are most and least likely to address relative to non-peripatetic users are depicted for each category in Figure 6. In both racist and misogynistic subreddits, peripatetic users are more involved in discourse about “social justice warriors” (discourse that is associated with both racism and misogyny (Brock 2015)). Additionally, in misogynistic subreddits, topics represented by the term “Nazi” and “trans” are more likely to be used by peripatetic users. From qualitative inspection, the topic represented by the word “sex” in the anti-LGBTQ category seems to capture discourse from TERF subreddits. This topic is much less likely to be used by peripatetic users. This could be because the other subreddit categories are likely to express views in opposition to all branches of feminism, including the type of feminism with which TERFs identify.
RQ3: Joining additional hate communities is associated with users become more active in hate groups and developing a broader hate-group lexicon. Complementing these results, Figure 4C shows the odds ratio of ingroup language after users join other hate subreddits (either of the same or a different category), and Figure 4D shows the percent activity increase. These results show that peripatetic users have a broader hate group lexicon, and become more active in hate groups overall. We caution that these results are not causal, as we do not compare against matched users who do not join additional hate subreddits. These results nonetheless motivate a root concern: not only is joining one hate subreddit harmful (Schmitz, Burghardt, and Muric 2022; Schmitz et al. 2024), but joining additional subreddits presents a compounding harm to users and social media.
@@
Racist .60 $\pm$ .002 .59 $\pm$ .002 .62 $\pm$ .002
Anti-LGBTQ .62 $\pm$ .003 .59 $\pm$ .004 .58 $\pm$ .004
Misogynistic .62 $\pm$ .003 .61 $\pm$ .003 .59 $\pm$ .003
@@
Table 1 lists the performance for our model predicting whether a user will post in a given category of subreddit. “Target” refers to posts made by the user for whom the prediction is being made, while “context” refers to posts to which they reply. The performance of each category excludes predictions for users whose initial posts were in that category. Uncertainties represent standard errors. For anti-LGBTQ and misogynistic subreddits, the model performs best when supplied with text written by the users and the text to which they replied, compared to either input type in isolation. However, the context alone is more predictive for users who originate in racist subreddits than both the context and the target.

### v2

@@ Methods
To understand whether users who initially post in hate subreddits are at risk of becoming engaged in hate subreddits of other types, we employ a process to match each user from a given hate subreddit (a “joiner”) with a similar “counterpart” user who did not engage in a hate subreddit before the joiner, as described in Schmitz et al. 2024. While we cannot answer whether joining one hate subreddit will lead a user to join a different type of hate subreddit, matching members of hate subreddits to users who are similar provides a strong baseline for assessing whether there is substantial overlap among the user bases of hate subreddits. Our matching process is summarized as follows: For each hate subreddit, we find non-hate subreddits in which users from the given hate subreddit are most active relative to the total size of the subreddit and then sample potential counterparts from those subreddits. Each user from the hate subreddit is matched to a counterpart using Mahalanobis distance matching. For a given hate subreddit, we only consider users who have posted in the subreddit as their first hate subreddit. The features used to measure the Mahalanobis distance between users are each user’s total Reddit “karma” (the number of total upvotes minus the number of total downvotes they receive on their posts), their total number of submissions, their total number of comments, their account creation date, and the total number of posts they made in the top 50 non-hate subreddits in which the members of the given hate subreddit were most active. For each counterpart user, all features were measured before the month the joiner posted in the hate subreddit. From these sets of users, we calculate the proportion who post in subreddits of a different type than the given hate subreddit that are joined within our chosen time threshold of six weeks following the users’ first post in the given subreddit. We report the proportions for each category of subreddit as a whole. As counterpart users by definition do not participate in the hate subreddit of their treatment counterpart, we record whether or not they become active in a different type of hate subreddit within the same time period as the hate subreddit user to whom they are matched.
In addition to quantifying the level of movement across subreddits through sets of matched users, we also calculate how likely users are to move to a certain category of subreddit given their original category. For each category, we calculate the proportion of peripatetic users who eventually move to each alternate category divided by a preferential attachment null model (Barabási and Albert 1999) in which peripatetic users join a category proportional to its number of users. This process provides us with a ratio for each pair of origin and destination subreddits, where a value greater than one indicates an unexpectedly greater influx of users from the origin subreddit category, and a value less than one indicates an unexpectedly smaller influx.
Peripatetic users and ingroup language
@@
Using the lexicon of ingroup language associated with each category of hate subreddit, we seek to understand how users who post in multiple categories of hate subreddits introduce such language to different types of communities. We identify the hate subreddit category each user first posts and the category they will eventually post in within six weeks. We then compare the language employed by peripatetic users in the first three days and prior to posting in other hate subreddits to the language used by those who exclusively posted within a single category in the first three days. This threshold captures users’ initial post behavior after adopting the initial subreddit but long before moving to a new hate subreddit category. To check robustness, we also collected all posts from users after joining the first hate subreddit but before posting in any other hate subreddit category. All results from these variations are qualitatively similar, as shown in the Appendix.
For each combination of origin and destination subreddits, we count the number of terms used from the lexicon of the destination category relative to the number of terms used from all other categories. These are counted separately for peripatetic users and non-peripatetic users. By counting the number of terms associated with each category, we are able to calculate the odds ratio that a peripatetic user employs terms from their destination category within the original category before they posted in the destination category, relative to non-peripatetic users. We assess the statistical significance of differences in language use between users who will eventually become peripatetic and those who will not via Fisher’s exact test. As there are some pairs of subreddits with relatively few peripatetic users, which can result in noisy odds ratios, we only report odds ratios for pairs of subreddits in which there are more than 20 peripatetic users that move from the origin subreddit to the destination subreddit.
We also measure the relative increase in ingroup language use from an alternate category after peripatetic users become active in said alternate category. For each peripatetic user, a symmetric window of time is used to measure behavior before and after becoming peripatetic (e.g., if a user posts in a misogynistic subreddit three weeks after posting in a racist subreddit, their behavior in that three weeks is measured as the “before” period and their behavior three weeks after posting in the misogynistic subreddit is measured as the “after” period).
@@
In addition to understanding how content associated with hate subreddits in which peripatetic users eventually post appears in their postings in their initial community, we are also interested in the general themes expressed by peripatetic users, unbounded by a specific source. To understand these themes, we employ topic modeling using BERTopic (Grootendorst 2022). BERTopic uses pre-trained transformer models to embed text, then clusters those embeddings to form topics. The topics are then summarized using the highest TF-IDF words for each cluster. For each hate subreddit, we fit a topic model using the first three days of posts for each user and prior to joining another subreddit category. As a preprocessing step, we remove URLs and the strings “[deleted]” and “[removed]” from all posts. Posts containing only URLs or these strings were removed from the dataset entirely. We use the “outlier reduction” feature of BERTopic to assign posts that were not assigned to a topic when initially fitting the model to the closest topic based on the cosine similarity of the text embeddings. We then merge the topic models of all subreddits of the same category. After merging the topic models, we count the number of posts in the top 100 topics that peripatetic users and non-peripatetic users employ. For each topic, the odds of a peripatetic user making a post that fits within the given topic are calculated relative to the odds of a non-peripatetic user making a post on the same topic. When comparing the odds ratios of topics, only topics that appear in at least ten percent of the subreddits within the given category are considered, as some topics may come from individual subreddits that are more or less likely to contain peripatetic users.
@@
To probe the dynamics wherein users join additional categories of hate subreddits, we initially seek to establish whether certain factors are predictive of a user joining a different category of hate subreddit in the first place. Specifically, as we establish how language spreads among the categories of hate subreddits, we are interested in how the text to which users are exposed, and the text that they themselves employ, is indicative of their participation in each type of hate subreddit. To examine the text to which users are exposed, we utilize the posts to which the user replies, as we are confident that the user has read these posts. However, it is important to note that this approach a) excludes a great deal of text that the users likely read within the given subreddit, and b) may reflect self-selection, as the text to which users reply is likely to be especially related to their interests. We frame this as a multi-label classification problem, where we build a neural network trained on the embeddings of each type of text, and have the neural network predict whether a user subsequently joins a subreddit in each category. The ground truth labels are thus binary labels indicating whether a given user posted in each subreddit category within six weeks of participating in their first hate subreddit. The text each user writes or replies to in any subreddit within three days of posting in their initial hate subreddit (including their first post in that subreddit) and prior to joining a new hate subreddit is embedded and fed into our model as input features (we explain embedding below).
As we are primarily interested in users posting in other categories of subreddits, and not continuing to post in their original category, we do not consider a user making multiple posts in their original category as a positive example. Using this strategy, we compile a dataset of approximately 100k randomly sampled users, 70% of which was used for training, with 15% being used each for validation and testing. As some categories (ableist and antisemitic) have very few examples of peripatetic users, we exclude them from the prediction model. In total, the model therefore predicts six categories.
@@
For each user in the dataset, all of the text that they post in their initial subreddit category is chronologically concatenated into the same document, with each post separated by the [SEP] token. The text to which the users reply is separately concatenated in the same way. Each document is embedded into a 768-dimensional vector using Longformer (Beltagy, Peters, and Cohan 2020). The vectors are then concatenated, resulting in a 1,538-dimensional input layer for the neural network. As some data are missing from Pushshift (Gaffney and Matias 2018), there are some cases in which all of the posts to which a user replied could not be retrieved; in those cases, a single token (‘UNK’) was used as a placeholder for the parent text. The category of subreddit in which each user initially posted, as well as the categories in which the users to whom they replied posted prior to their interaction, are input as a one-hot encoding vector into the neural network at the second layer. The final neural network has three layers, with 780 dimensions in the second layer (where 12 of the input features derive from the one-hot encoding describing the initial subreddit and the subreddits of the parent users) and 384 dimensions in the third layer (i.e., each intermediate layer divides the number of features from the text vectors by a factor of two, with additional features in the second layer due to the one-hot encoding). At each hidden layer, a dropout of 0.6 is used, as well as the leaky ReLU activation function. The sigmoid activation function is used at the final output layer, and the model is fit with the binary cross-entropy loss function. We perform the training on an internal computing cluster using an NVIDIA RTX A6000 GPU.
@@
Finally, as there is variation in the performance of the model among individual runs, we take random splits of training, validation, and testing. This entails using a different seed for the Python library, sklearn’s train_test_split function. After each new split, we train the model to calculate the ROC-AUC. This is performed 50 times to determine the mean and standard errors of the ROC-AUC. In addition, we only report performance for each category considering users who did not originate in that category, as we do not consider movement within the same category.
@@ Results
Figure 2 shows the average number of posts predicted as hate speech for each cluster identified by the $k$-means algorithm. From the distributions, it is clear that the clusters mostly align with individual bias categories, such as race or gender. However, some categories exhibit a combination of bias categories: whereas the “general hate” category is not particularly high in any bias category, the “anti-LGBTQ” category is high in both homophobia and transphobia. With the exception of these categories, we refer to each category of subreddits as the highest-scoring bias category in each subreddit. The counts of subreddits and users per cluster are displayed in Table 2. Figure 1 shows the distributions of average hate scores for each subreddit plotted in two-dimensional space. Observing the plot, the categories form distinct clusters.
Upon qualitative inspection of the subreddits within each cluster, they seem to thematically fit with the types of hate predicted as most prevalent within each cluster. For example, the anti-LGBTQ cluster contains TERF subreddits such as r/GenderCritical or homophobic subreddits such as r/homophobes. The general hate cluster contains edgy humor subreddits such as r/CringeAnarchy and r/ImGoingToHellForThis, and the xenophobic cluster contains subreddits such as r/The_Donald which express strong anti-immigrant sentiment. The racist cluster contains subreddits such as r/GreatApes and r/CoonTown, in which anti-Black racism is frequent.
@@
Figure 3: Proportion of users who post in alternative subreddit categories within six weeks after becoming active in their initial hate subreddit category. Error bars represent standard errors.
@@
Figure 3 displays the proportion of newcomers who post in hate subreddits with matched treatment/counterpart pairs. It is apparent that, on average, the users who post in each hate subreddit subsequently participate in many more subreddits of another type than do users who do not post in the given hate subreddit (two times higher depending on the type of subreddit; unweighted mean across all eight categories: 2.1). Using the two-proportion Z test to assess differences between treatment and counterpart groups for each category, we find that the observed results are extremely unlikely to happen by chance, with p-values $<10^{-5}$ for each result.
RQ1: Categories of hate subreddits show preferences for other categories.
@@
Figure 4: Heatmaps representing peripatetic user behavior (A) Rate of users from the original subreddit type that subsequently posted in the alternate subreddit. (B) Odds ratios of peripatetic vs. non-peripatetic users using language from the alternate hate subreddit lexicon within the origin subreddit type. (C) Change in use of language from alternate hate subreddit lexicons after posting in alternate hate subreddits. In (B) and (C), cells with numbers in them represent statistical significance (p-values $<0.05$).
Figure 4A displays the rate at which members who originate in each subreddit category subsequently post in other categories of subreddits. From the figure, it is apparent that, for most categories, users are more likely than chance to participate again in the category they originated in. Additionally, antisemitic, racist, Islamophobic, and xenophobic subreddits exhibit higher user overlap with each other than expected by chance. Users in the ableist category show a preference for general hate subreddits and vice versa.
@@
RQ2: Users who will become peripatetic introduce language associated with alternate hate subreddits to the hate subreddit category in which they initially posted.
@@
Figure 4B displays the odds of peripatetic users posting material, in the category to which they first contributed, that employs language common in the peripatetic user’s subsequent subreddit category, relative to the odds that non-peripatetic users employ such language in the first category. The heatmap reveals that, for the most part, the results are statistically significant and are associated with peripatetic users employing more ingroup language associated with the destination subreddit relative to non-peripatetic users. While there are some pairs of categories where we observe the opposite result, these tend not to be statistically significant.
@@
Figure 5 displays the ten most common topics within each category of subreddit. Among these topics, it is apparent that many of the topics are related to the bias categories identified by our hate speech detection model. For example, the top topic in the antisemitic category is represented by the word “holocaust.” In the misogynistic category, discussions about feminism, and perceived attractiveness, and topics represented by misogynistic slurs are among the most frequent. In contrast, the general hate category contains discussion about various identities, including topics related to race and feminism.
Figure 6: Topics with the five highest/lowest log-odds ratios of use between peripatetic users and non-peripatetic users. Positive values indicate topics that are over-represented among peripatetic users. Topics with an adjusted p-value greater than 0.05 according to Fisher’s exact test are lighter to show statistical insignificance. Stopwords were removed from topic representations.
The topics that peripatetic users are most and least likely to address relative to non-peripatetic users are depicted for each category in Figure 6. In general hate subreddits, topics represented by misogynistic slurs, and those discussing incels or relationships are most highly associated with peripatetic activity, indicating that the users discussing these topics may later move to manosphere subreddits in the misogynistic category. Meanwhile, the topics most positively associated with peripatetic haters in the racist subreddits are comparatively more extreme than those more likely to be employed by non-peripatetic users. In the Islamophobic category, topics represented by misogynistic slurs and the word “girl” are more likely to be used by peripatetic users.
RQ3: Joining additional hate communities is associated with users developing a broader hate-group lexicon. Complementing these results, Figure 4C shows the odds ratio of ingroup language after users join other hate subreddits (either of the same or a different category). These results show that peripatetic users have a broader hate group lexicon. We caution that this result is not causal, as we do not compare against matched users who do not join additional hate subreddits. These results nonetheless motivate a root concern: not only is joining one hate subreddit harmful (Schmitz, Burghardt, and Muric 2022; Schmitz et al. 2024), but joining additional subreddits presents compounding harm to users and social media. In addition to exploring the change in language use of peripatetic users, we also considered their changes in activity rates. However, we found no significant difference in activity rates before and after users became peripatetic, nor was the change in peripatetic users’ activity substantially different from the activity of non-peripatetic users who continued posting in their original category.
@@
General hate .62 +/- .002 .60 +/- .002 .58 +/- .002
Misogynistic .61 +/- .002 .60 +/- .003 .56 +/- .003
Anti-LGBTQ .63 +/- .009 .63 +/- .009 .58 +/- .009
Racist .68 +/- .009 .66 +/- .009 .60 +/- .010
Xenophobic .67 +/- .002 .65 +/- .002 .58 +/- .003
Islamophobic .67 +/- .006 .64 +/- .006 .62 +/- .006
@@
Table 1 lists the performance for our model predicting whether a user will post in a given category of subreddit distinct from their initial category. “Target” refers to posts made by the user for whom the prediction is being made, while “context” refers to posts to which they reply. Uncertainties represent standard errors. For almost all categories of subreddits, the model performs best when supplied with text written by the users and the text to which they replied, compared to either input type in isolation. Notably, the text and metadata of posts users write is more predictive of peripatetic activity than the text and metadata of the users to whom they reply.